# BroadGen: A Framework for Generating Effective and Efficient Advertiser Broad Match Keyphrase Recommendations

### v2

@@ 3.1. Input Generation
Two input matrices are constructed, Anchor Similarity Matrix $A^{|Q|\times|t|}$ and Query-Query Similarity Matrix $S^{|Q|\times|Q|}$, where ||$\cdot$|| ⋅ | determines cardinality. In constructing $A$, a $|t|$-dimensional vector is generated for each query. This vector comprises of binary values, where 1s and 0s indicate the query token’s presence in the title. The construction of $S$ is done by computing the common token count between each pair of queries in $Q$. Each row of $S$ can be considered as a $|Q|$-dimensional vector that represents each query depending on similarity with other queries. Each vector of matrix $S$ is normalized by dividing by its maximum, while each vector of the matrix $A$ is normalized by its sum. The Query-Query Similarity matrix $S$ is transformed into a distance matrix by $D_{ij}=1-S_{ij}$.
@@
9:        $C\leftarrow pick(C,|K|)$triangleright Pick largest clusters.
@@ 3.2. Clustering Core
The core functionality starts with computing the pairwise distance between each pair of vectors $z_{i},z_{j}\in\hat{Z}$ as demonstrated in Equation 2. Now, with the pairwise distance matrix $\tilde{Z}$, the core performs agglomerative clustering (Müllner, 2011) by constructing linkage distances with ward variance minimization (Jr., 1963) and optimal leaf ordering (Bar-Joseph et al., 2001). We denote the clustering function as $R(\cdot,\cdot)$.
(2) $\tilde{Z}=1-\frac{z_{i}\cdot z_{j}}{||z_{i}||_{2}\cdot||z_{j}||_{2}}$
$R(\tilde{Z},\tilde{r})$ outputs a set of clusters $C={c_{1},c_{2},...}$, where each cluster $c_{i}$ groups similar queries. While, $\tilde{r}$ is a crucial hyper-parameter called as the threshold ratio that determines how many clusters are formed. [^3] We use the inconsistent criterion that compares the linkage distance of current and previous levels of the hierarchy. A fixed value of $\tilde{r}$ can lead to more or less number of clusters $|C|$ than the requirement $|K|$. The Algorithm 1 shows how the required number of clusters is retrieved, through decreasing the threshold by epsilon (line 6) when there are less clusters. If there are more clusters than required, it picks the top $|K|$ clusters (line 9) in decreasing order of the cluster size.
3.3. Representation Generation
@@
A representation for each cluster in $C$ depicts a keyphrase that can be extended to queries in the cluster. In the final core, the representation is determined as the set of tokens common among the tokenized queries within each cluster. At times, clusters can be large leading to single or no tokens (in absence of commonality). In these situations, an optimal subset of the cluster is identified whose representation is more than one token. The common tokens from each cluster need to be ordered before concatenating into a string. We employ a graph-based approach to infer an ordering of the common tokens based on their relative positions across the queries. A graph is constructed with all tokens as the vertices; for each co-occurring pair of tokens a directed edge is added indicating local precedence. A topological sort (Tarjan, 1976) of the graph using a depth-first search algorithm produces a consistent global ordering of the tokens. This method of determining precedence is rooted in prior work, such as sentence ordering task (Prabhumoye et al., 2020). For each cluster, the common tokens are arranged following the global order and represented as a string.
Figure 4. Illustration of BroadGen.
3.4. Implementation Details
@@
We show an example of our model on three different items in Figure 4. For each item, only one cluster of queries is shown along with the keyphrases that are generated.  [^4] This is just for illustration, there can be multiple clusters per item depending on the number of input queries. Aligning with our goals described in Section 1.2.4, we develop the framework in Python to enable its integration into eBay’s in-house Spark system. This off-loads the overhead of data transfer and parallelism to the spark system, making it feasible to scale for billions of items while mitigating engineering costs. The cores of the framework are designed to be fast and efficient. The asymptotic time complexity of all the cores is $O(|Q|^{2})$ which might seem large but $Q$ is per item which is typically in the thousands.
@@
The framework’s recommendations are adaptive to the input queries — depending on number and quality of queries, the recommendations change significantly. Clustering the queries helps us optimize the reach of a particular keyphrase by limiting it’s tokens to emerge from the cluster which assists to some extent in identifying relevant reach (to be discussed later in Section 4.2.2). To obtain the cluster representation, we impose specific constraints to prevent the inclusion of generic broad keyphrases from the cluster. If the retrieved representation consists of a single token, we instead use a subset of the cluster for retrieval. The subset is chosen from all possible combinations of size $|C|-1$ where the retrieved representation is more than a single token.
4. Experimentation
4.1. Setup
4.1.1. OOT Formulation
@@
To evaluate the effectiveness of broad match keyphrases in addressing the challenge of shifting label distribution, we assess their generalized performance in an Out-of-Time (OOT) (Zhang et al., 2025) setting.  [^5] Commonly referred to as out-of-distribution generalization (Li et al., 2023; Yu et al., 2024; Liu et al., 2025). Specifically, the algorithms for generating broad match keyphrases use as input the historical queries searched by buyers prior to time $T$, termed as Pre-queries and evaluated against the query distribution observed in the subsequent $T+30$ day period called as Post-queries.
@@
Table 1. Details of meta categories from eBay.
4.1.2. Data Augmentation
The primary limitation of BroadGen’s methodology lies in its dependency on historical interaction data between items and queries for the generation of keyphrases. Although the algorithm functions effectively under warm start scenarios, it encounters difficulties with new or “cold” items that have limited interaction data available. To address this shortcoming, our approach incorporates input from state-of-the-art models deployed at eBay to provide precise keyphrase recommendations for cold-start items (Mishra et al., 2024, 2025), thereby bolstering the performance of our algorithm when dealing with cold-start conditions. [^6] This kind of data augmentation has been studied in (Zhang et al., 2025).
4.1.3. Datasets
We present results on 4 meta-categories of eBay, CAT_1 to CAT_4, each representing a group whose size is based on number of items and queries. The category sizes are largest from CAT_1 to CAT_4, the detailed statistics for each category is presented in Table 1. For each category, each item’s future query distribution $Q$ is limited to 1000 queries. Each of these 1000 queries has at least 1 token overlap with the item’s title.
4.2. Evaluation Framework
4.2.1. Defining Relevance
In eBay’s Advertising ecosystem, a downstream relevance model filters the queries matched by broad match keyphrases, which is done by eBay Search. Hence, it is crucial to avoid unnecessary recommendations that are too generic which impacts the downstream efficiency and focus on passing the subsequent filtering processes by Search on the matched queries. In addition, the queries have to be relevant to the item in terms of human judgment, or else it will not be engaged by the buyers. To assess the relevance between item and query, we employ a BERT (Devlin et al., 2019) cross-encoder model (Wolf et al., 2020) that has been fine-tuned using judgment data derived from Large Language Models (LLMs) at eBay (Dey et al., 2025a). This model’s relevance judgment emulates human judgment while still aligning with eBay Search’s definition of relevance. We use this relevance judgment to check whether a query is relevant to an item for our evaluation. Employing this cross-encoder’s judgment also captures the semantic nuances of relevance that traditional lexical scoring methods generally overlook (Dey et al., 2025b).
4.2.2. Relevant Reach Estimation (RRE)
@@
PreQuery Model 0.478 0.517 0.498 0.633 0.319 0.378 0.315 0.337 0.383 0.437 0.386 0.440
Transfusion 0.535 0.558 0.527 0.628 0.194 0.213 0.181 0.203 0.285 0.308 0.270 0.382
BroadGen 0.474 0.534 0.543 0.716 0.145 0.296 0.209 0.371 0.223 0.381 0.302 0.489
BroadGen(Aug) 0.492 0.510 0.503 0.723 0.375 0.538 0.438 0.625 0.426 0.524 0.468 0.671
Oracle 0.320 0.460 0.440 0.780 1.000 1.000 1.000 1.000 0.480 0.630 0.610 0.880
Table 2. Precision, recall and F1 metrics for all the models. Darker hues represent more favorable outcomes.
The relevant reach of a broad match keyphrase is subject to the distribution of buyers’ search queries, the broad matching mechanism and sellers’ advertising bids and budget. Although it’s infeasible to compute this metric offline, we can approximate it through sampling from the Post-query distribution. In addition, we define a relevance function $\hat{R}(i,q)$ based on the definition relevance explicated in 4.2.1 and a broad match function B(k, q) as followed:
(3) $\hat{R}(i,q)=\begin{cases}1&\text{if $q$ is relevant to $i$}\\ 0&\text{if $q$ is not relevant to $i$}\end{cases}$
(4) $B(k,q)=\begin{cases}1&\text{$k$ can be broad matched to $q$}\\ 0&\text{$k$ cannot be broad matched to $q$}\end{cases}\$
(5) $B@\mathcal{K}(k,q)=\mathds{1}_{\sum_{k\in\mathcal{K}}B(k,q)>0}\$
@@
(6) $Precision@\mathcal{K}(i,k)=\frac{\sum\limits_{q\sim Q}B@\mathcal{K}(k,q)\cdot\hat{R}(i,q)}{\sum\limits_{q\sim Q}B@K(k,q)}$
@@
In practice, we choose the broad match criterion for $B(k,q)$ to be 1 if the tokens of $k$ are a subset of the tokens of $q$. We relax this criterion when $k$ contains N tokens ($N>=3$) where we require $q$ to contain at least $N-1$ tokens from $k$.
4.2.3. Proportional Token Reach (PTR)
@@
(7) $PTR(k,q)=\frac{|q\bigcap k|}{|q\bigcap k|+\alpha\cdot|q-k|+\beta\cdot|k-q|}$
@@
4.3. Offline Performance
We compare with the Transfusion and the Match Type Curation models from Section 2. As a proxy for the match type curation strategy, we use the top pre-queries (used as input for BroadGen) as the recommended broad match type keyphrases and term it as the Prequery model. We generate 5 broad match type keyphrases for each model. BroadGen(Aug) indicates BroadGen with the data augmentation.
4.3.1. RRE Results
We show here the RRE’s precision and recall scores as defined in Section 4.2.2 relative to the prequery model on the 4 categories of eBay in Table 2. We use the traditional definition of F1 on top our precision and recall scores from Equation 6. BroadGen(Aug) achieves the highest recall and F1 scores, despite having precision similar to the pre-query model. While generally exhibiting high precision, Transfusion has the lowest recall and F1 scores. Intuitively, recall represents the effectiveness of broad match keyphrases by measuring their ability to capture a comprehensive set of queries relevant to the seed item. Precision, on the other hand, reflects the accuracy of these broad keyphrases, indicating the overall relevance of the buyer queries matched by Search.
To better understand how does RRE fairs when there is a perfect baseline, we design an Oracle that reports all the 1000 queries for any item as broad matched ($B(\cdot)=1$), which includes both BERT relevant and non-relevant queries leading to a 100% recall model. Although the recall of this model is 100%, as seen in Table 2, the precision is generally the lowest. This is due to the large number of queries the Oracle covers while the numerator is limited to the BERT relevant queries for each item. In essence, not every query reachable by a keyphrase is relevant to the item thus a balance between the precision and recall is essential.
@@ 4.3.1. RRE Results
BroadGen (Aug) 1.0 1.5 1.16$\times$ 1.08$\times$ 1.18$\times$ 1.19$\times$
1.5 1.0 1.08$\times$ 1.01$\times$ 1.12$\times$ 1.12$\times$
Transfusion 1.0 1.5 0.64$\times$ 0.71$\times$ 0.43$\times$ 0.55$\times$
1.5 1.0 0.61$\times$ 0.67$\times$ 0.40$\times$ 0.52$\times$
Table 3. Average PTR relative to the prequery model. Darker hues show more favorable outcomes.
4.3.2. PTR Results
Table 3 shows the average PTR scores as defined in Section 4.2.3 of the BroadGen(Aug) model relative to the Prequery model. A higher PTR score indicates how the keyphrases lexically resemble the queries — it takes few modifications to a keyphrase to match a relevant query. The observations are shown with two values of $\alpha$ and $\beta$ that control the penalty for dissimilarity. For both values of $\alpha$ and $\beta$, BroadGen(Aug) performs better in all categories. When $\alpha=1.0$ and $\beta=1.5$, PTR penalizes extra words in the keyphrases more than missing words in the queries and when $\alpha=1.5,\beta=1.0$ it does vice-versa. Our model performs relatively better in the first case indicating that it requires less deletion of tokens to match the query. This is crucial as many platform’s implementation of broad matching prefer not to delete tokens. BroadGen and Prequery models’ similar performance in the second case suggests that our recommendations are not too generic. Transfusion being generative and not informed by any historical precedence, many of its keyphrases’ tokens do not match that of the queries resulting in lower PTR.
Figure 5. Comparison of Precision, Recall and F1 for BroadGen(Aug) across keyword counts and categories.
4.4. Ablation Studies
4.4.1. Dependency on Input Queries
The performance of BroadGen(Aug) in contrast to BroadGen in Table 2 indicates the importance of data augmentation for the quality of recommendations. Across all categories, recall and F1 scores experience a sharp decline, while precision changes minimally. Although precision improves for CAT_2 and CAT_3, it does not significantly compensate for the drop in recall.
To enhance our understanding, we examine the percentage change in BroadGen’s scores when relying solely on pre-queries. Table 4 illustrates a notable decline in recall and F1 scores for items with fewer than 10 input pre-queries. This situation is akin to dealing with cold items, where enough related queries are not available. With a limited number of input queries, there is insufficient data to effectively cluster and generate keyphrases. The improved performance of items with more than 10 pre-queries highlights the importance of increasing the number of queries, aided by recommendations from other recall models, which proved to be transformative for BroadGen(Aug).
Cats. # Pre-queries ¿= 10 # Pre-queries ¡ 10

### v3

@@ III-B Input Generation
Two input matrices are constructed, Anchor Similarity Matrix $A^{|Q|\times|t|}$ and Query-Query Similarity Matrix $S^{|Q|\times|Q|}$, where $|\cdot|$ determines cardinality. In constructing $A$, a $|t|$-dimensional vector is generated for each query. This vector comprises of binary values, where 1s and 0s indicate the query token’s presence in the title. The construction of $S$ is done by computing the common token count between each pair of queries in $Q$. Each row of $S$ can be considered as a $|Q|$-dimensional vector that represents each query depending on similarity with other queries. Each vector of matrix $S$ is normalized by dividing by its maximum, while each vector of the matrix $A$ is normalized by its sum. The Query-Query Similarity matrix $S$ is transformed into a distance matrix by $D_{ij}=1-S_{ij}$.
@@
9:   $C\leftarrow pick(C,|K|)$ $\triangleright$ Pick largest clusters.
@@ III-C Clustering Core
The core functionality starts with computing the pairwise distance between each pair of vectors $z_{i},z_{j}\in\hat{Z}$ as demonstrated in Equation 2. Now, with the pairwise distance matrix $\tilde{Z}$, the core performs agglomerative clustering [36] by constructing linkage distances with ward variance minimization [37] and optimal leaf ordering [38]. We denote the clustering function as $R(\cdot,\cdot)$.
[…]
III-D Representation Generation
@@
A representation for each cluster in $C$ depicts a keyphrase that can be extended to queries in the cluster. In the final core, the representation is determined as the set of tokens common among the tokenized queries within each cluster. At times, clusters can be large leading to single or no tokens (in absence of commonality). In these situations, an optimal subset of the cluster is identified whose representation is more than one token. The common tokens from each cluster need to be ordered before concatenating into a string. We employ a graph-based approach to infer an ordering of the common tokens based on their relative positions across the queries. A graph is constructed with all tokens as the vertices; for each co-occurring pair of tokens a directed edge is added indicating local precedence. A topological sort [39] of the graph using a depth-first search algorithm produces a consistent global ordering of the tokens. This method of determining precedence is rooted in prior work, such as sentence ordering task [40]. For each cluster, the common tokens are arranged following the global order and represented as a string.
Figure 4: Illustration of BroadGen.
III-E Implementation Details
@@
We show an example of our model on three different items in Figure 4. For each item, only one cluster of queries is shown along with the keyphrases that are generated.  [^4] This is just for illustration, there can be multiple clusters per item depending on the number of input queries. Aligning with our goals described in Section I-B4, we develop the framework in Python to enable its integration into eBay’s in-house Spark system. This off-loads the overhead of data transfer and parallelism to the spark system, making it feasible to scale for billions of items while mitigating engineering costs. The cores of the framework are designed to be fast and efficient. The asymptotic time complexity of all the cores is $O(|Q|^{2})$ which might seem large but $Q$ is per item which is typically in the thousands.
@@
The framework’s recommendations are adaptive to the input queries — depending on number and quality of queries, the recommendations change significantly. Clustering the queries helps us optimize the reach of a particular keyphrase by limiting it’s tokens to emerge from the cluster which assists to some extent in identifying relevant reach (to be discussed later in Section IV-B2). To obtain the cluster representation, we impose specific constraints to prevent the inclusion of generic broad keyphrases from the cluster. If the retrieved representation consists of a single token, we instead use a subset of the cluster for retrieval. The subset is chosen from all possible combinations of size |C|1|C|-1 where the retrieved representation is more than a single token.
III-F Explainability
Since BroadGen is based on token correspondence between candidate exact match keyphrases (queries) with an underlying pre-query population coming from historical buyer interaction, it is explainable. Given a BroadGen keyphrase we would be able to provide the exact clustering of the input pre-queries that were considered and the token distribution within the cluster, summarizing the differences between the queries as illustrated in Figure 4.
IV Experimentation
IV-A Setup
IV-A1 OOT Formulation
@@
To evaluate the effectiveness of broad match keyphrases in addressing the challenge of shifting label distribution, we assess their generalized performance in an Out-of-Time (OOT) [20] setting.  [^5] Commonly referred to as out-of-distribution generalization [41, 42, 43]. Specifically, the algorithms for generating broad match keyphrases use as input the historical queries searched by buyers prior to time $T$, termed as Pre-queries and evaluated against the query distribution observed in the subsequent $T+30$ day period called as Post-queries.
IV-A2 Data Augmentation
The primary limitation of BroadGen’s methodology lies in its dependency on historical interaction data between items and queries for the generation of keyphrases. Although the algorithm functions effectively under warm start scenarios, it encounters difficulties with new or “cold” items that have limited interaction data available. To address this shortcoming, our approach incorporates input from state-of-the-art models deployed at eBay to provide precise keyphrase recommendations for cold-start items [3, 2], thereby bolstering the performance of our algorithm when dealing with cold-start conditions. [^6] This kind of data augmentation has been studied in [20].
@@ IV-A2 Data Augmentation
TABLE I: Details of meta categories from eBay.
IV-A3 Datasets
We present results on 4 meta-categories of eBay, CAT_1 to CAT_4, each representing a group whose size is based on number of items and queries. The category sizes are largest from CAT_1 to CAT_4, the detailed statistics for each category is presented in Table I. For each category, each item’s future query distribution $Q$ is limited to 1000 queries. Each of these 1000 queries has at least 1 token overlap with the item’s title.
@@ IV-A3 Datasets
PreQuery Model 0.48 0.52 0.50 0.63 0.32 0.38 0.32 0.34 0.38 0.44 0.39 0.44
Transfusion 0.54 0.56 0.53 0.63 0.19 0.21 0.18 0.20 0.29 0.31 0.27 0.38
BroadGen 0.47 0.53 0.54 0.72 0.15 0.30 0.21 0.37 0.22 0.38 0.30 0.49
BroadGen(Aug) 0.49 0.51 0.50 0.72 0.38 0.54 0.44 0.63 0.43 0.52 0.47 0.67
Oracle 0.32 0.46 0.44 0.78 1.00 1.00 1.00 1.00 0.48 0.63 0.61 0.88
TABLE II: Precision, recall and F1 metrics for all the models. Darker hues represent more favorable outcomes.
IV-B Evaluation Framework
IV-B1 Defining Relevance
In eBay’s Advertising ecosystem, a downstream relevance model filters the queries matched by broad match keyphrases, which is done by eBay Search. Hence, it is crucial to avoid unnecessary recommendations that are too generic which impacts the downstream efficiency and focus on passing the subsequent filtering processes by Search on the matched queries. In addition, the queries have to be relevant to the item in terms of human judgment, or else it will not be engaged by the buyers. To assess the relevance between item and query, we employ a BERT [44] cross-encoder model [45] that has been fine-tuned using judgment data derived from Large Language Models (LLMs) at eBay [12]. This model’s relevance judgment emulates human judgment while still aligning with eBay Search’s definition of relevance. We use this relevance judgment to check whether a query is relevant to an item for our evaluation. Employing this cross-encoder’s judgment also captures the semantic nuances of relevance that traditional lexical scoring methods generally overlook [8].
IV-B2 Relevant Reach Estimation (RRE)
The relevant reach of a broad match keyphrase is subject to the distribution of buyers’ search queries, the broad matching mechanism and sellers’ advertising bids and budget. Although it’s infeasible to compute this metric offline, we can approximate it through sampling from the Post-query distribution. In addition, we define a relevance function $\hat{R}(i,q)$ based on the definition relevance explicated in IV-B1 and a broad match function B(k, q) as followed:
[…]
$B@\mathcal{K}(k,q)=\mathds{1}_{\sum_{k\in\mathcal{K}}B(k,q)>0}\\$ (5)
@@ IV-B2 Relevant Reach Estimation (RRE)
$Precision@\mathcal{K}(i,k)=\frac{\sum\limits_{q\sim Q}B@\mathcal{K}(k,q)\cdot\hat{R}(i,q)}{\sum\limits_{q\sim Q}B@K(k,q)}$ (6)
@@
In practice, we choose the broad match criterion for $B(k,q)$ to be 1 if the tokens of $k$ are a subset of the tokens of $q$. We relax this criterion when $k$ contains N tokens ($N>=3$) where we require $q$ to contain at least N1N-1 tokens from $k$.
IV-B3 Proportional Token Reach (PTR)
@@
$PTR(k,q)=\frac{|q\bigcap k|}{|q\bigcap k|+\alpha\cdot|q-k|+\beta\cdot|k-q|}\vskip 2.84526pt$ (7)
@@
IV-C Offline Performance
We compare with the Transfusion and the Match Type Curation models from Section II. As a proxy for the match type curation strategy, we use the top pre-queries (used as input for BroadGen) as the recommended broad match type keyphrases and term it as the Prequery model. We generate 5 broad match type keyphrases for each model. BroadGen(Aug) indicates BroadGen with the data augmentation.
IV-C1 RRE Results
We show here the RRE’s precision and recall scores as defined in Section IV-B2 relative to the prequery model on the 4 categories of eBay in Table II. We use the traditional definition of F1 on top our precision and recall scores from Equation 6. BroadGen(Aug) achieves the highest recall and F1 scores, despite having precision similar to the pre-query model. While generally exhibiting high precision, Transfusion has the lowest recall and F1 scores. Intuitively, recall represents the effectiveness of broad match keyphrases by measuring their ability to capture a comprehensive set of queries relevant to the seed item. Precision, on the other hand, reflects the accuracy of these broad keyphrases, indicating the overall relevance of the buyer queries matched by Search.
To better understand how does RRE fairs when there is a perfect baseline, we design an Oracle that reports all the 1000 queries for any item as broad matched ($B(\cdot)=1$), which includes both BERT relevant and non-relevant queries leading to a 100% recall model. Although the recall of this model is 100%, as seen in Table II, the precision is generally the lowest. This is due to the large number of queries the Oracle covers while the numerator is limited to the BERT relevant queries for each item. In essence, not every query reachable by a keyphrase is relevant to the item thus a balance between the precision and recall is essential.
IV-C2 PTR Results
Table III shows the average PTR scores as defined in Section IV-B3 of the BroadGen(Aug) model relative to the Prequery model. A higher PTR score indicates how the keyphrases lexically resemble the queries — it takes few modifications to a keyphrase to match a relevant query. The observations are shown with two values of $\alpha$ and $\beta$ that control the penalty for dissimilarity. For both values of $\alpha$ and $\beta$, BroadGen(Aug) performs better in all categories. When $\alpha=1.0$ and $\beta=1.5$, PTR penalizes extra words in the keyphrases more than missing words in the queries and when $\alpha=1.5,\beta=1.0$ it does vice-versa. Our model performs relatively better in the first case indicating that it requires less deletion of tokens to match the query. This is crucial as many platform’s implementation of broad matching prefer not to delete tokens. BroadGen and Prequery models’ similar performance in the second case suggests that our recommendations are not too generic. Transfusion being generative and not informed by any historical precedence, many of its keyphrases’ tokens do not match that of the queries resulting in lower PTR.
@@ IV-C2 PTR Results
BroadGen
[…]
TABLE III: Average PTR relative to the prequery model. Darker hues show more favorable outcomes.
IV-D Ablation Studies
IV-D1 Dependency on Input Queries
The performance of BroadGen(Aug) in contrast to BroadGen in Table II indicates the importance of data augmentation for the quality of recommendations. Across all categories, recall and F1 scores experience a sharp decline, while precision changes minimally. Although precision improves for CAT_2 and CAT_3, it does not significantly compensate for the drop in recall.
To enhance our understanding, we examine the percentage change in BroadGen’s scores when relying solely on pre-queries. Table IV illustrates a notable decline in recall and F1 scores for items with fewer than 10 input pre-queries. This situation is akin to dealing with cold items, where enough related queries are not available. With a limited number of input queries, there is insufficient data to effectively cluster and generate keyphrases. The improved performance of items with more than 10 pre-queries highlights the importance of increasing the number of queries, aided by recommendations from other recall models, which proved to be transformative for BroadGen(Aug).
Cats. # Pre-queries >=>= 10 # Pre-queries << 10